OpenStudy (bella21121):

Approximately how many meters are equal to 11.4 ft? Note: 1 m ≅ 3.3 ft Round your answer to the nearest hundredths. 3.45 m 8.10 m 14.70 m 37.62 m

OpenStudy (kittiwitti1):

conversion factor: \(\Large{\frac{1m}{3.3ft}}\) on average. they want you to find the \(\color{red}{\text{value given in meters}}\) dimensional analysis:\[\large{11.4ft\times\frac{1m}{3.3ft}}\]

OpenStudy (kittiwitti1):

there is a similar unit in the numerator and denominator (ft):\[\large{11.4\cancel{\color{red}{ft}}\times\frac{1m}{3.3\cancel{\color{red}{ft}}}}\]this gets you to \(\Large{11.4\times\frac{1m}{3.3}}\) or -\[\huge{\frac{11.4\times1}{3.3}m}\]simplify this and you have the answer

OpenStudy (bella21121):

alright thank you i had a problem with answering this ^^'

OpenStudy (kittiwitti1):

no prob ☺ happy openstudying ♣

OpenStudy (bella21121):

^^ bye